OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 22 and the last term is 1,299,078?

OpenStudy (ashleyisakitty):

a =22 & ar^5=1299078 r^5=1299078/22=59049 r=9 S(6)=22(9^6-1)/9-1)=22(531441-1)/8=1461460

OpenStudy (anonymous):

you're awesome

OpenStudy (ashleyisakitty):

:3